# Hypergraph Foundation Model

### v1

@@ V-D3 Ablation Study on Structure Knowledge Extraction
Number of Clustering Number of Classes
1 2 3 4 5 6
@@
Movielens-Text $0.2884_{\pm.02}$ $0.3132_{\pm.09}$ $0.3281_{\pm.07}$ $0.3304_{\pm.05}$ $\mathbf{0.3464_{\pm.08}}$ $0.3217_{\pm.07}$ 5
GoodBook-Text $0.3378_{\pm.06}$ $0.3307_{\pm.03}$ $0.3294_{\pm.03}$ $0.3478_{\pm.05}$ $\mathbf{0.3588_{\pm.04}}$ $0.3227_{\pm.04}$ 4
PPI-Text $0.4919_{\pm.09}$ $0.5371_{\pm.09}$ $0.5581_{\pm.09}$ $\mathbf{0.5613_{\pm.09}}$ $0.5371_{\pm.08}$ $0.5016_{\pm.06}$ 3
@@ VI Conclusion
In this work, We introduce Hyper-FM, the first Hypergraph Foundation Model for multi-domain knowledge extraction. By implementing Hierarchical High-Order Neighbor Guided Vertex Knowledge Embedding, we integrate domain-specific information into vertex features, addressing varying dimensions and structural complexities. Our Hierarchical Multi-Hypergraph Guided Structural Knowledge Extraction method allows scalable and efficient extraction of structural knowledge from diverse datasets, mitigating negative transfer and domain-specific blending. Additionally, we curate 10 text-attributed hypergraph (TAHG) datasets to bridge hypergraph neural networks and large language models. Empirical evaluations show that Hyper-FM outperforms baseline methods by approximately 10% across these datasets, demonstrating its robustness and effectiveness. We also establish a novel scaling law, revealing that increasing domain diversity significantly enhances model performance, unlike merely expanding vertex and hyperedge counts. Hyper-FM sets a new benchmark for hypergraph foundation models, providing a scalable and effective solution for multi-domain knowledge extraction. Future work will explore expanding TAHG datasets and applying Hyper-FM to a broader range of domains to further enhance its versatility and impact.

### v2

@@ V-D2 Ablation Study on Hypergraph Sampling Strategy
Aminer-Text $0.4217_{\pm.02}$ $\mathbf{0.4325_{\pm.02}}$
Cora-CA-Text $0.2462_{\pm.02}$ $\mathbf{0.2633_{\pm.03}}$
Cora-$C$-Text $0.3224_{\pm.05}$ $\mathbf{0.3706_{\pm.04}}$
Pubmed-$C$-Text $0.3902_{\pm.01}$ $\mathbf{0.4009_{\pm.01}}$
Arxiv-Text $0.3960_{\pm.05}$ $\mathbf{0.4189_{\pm.04}}$
IMDB-Text $0.2661_{\pm.04}$ $\mathbf{0.2676_{\pm.04}}$
OGBN-Arxiv $0.2655_{\pm.02}$ $\mathbf{0.2789_{\pm.01}}$
@@ V-D3 Ablation Study on Structure Knowledge Extraction
Number of Clustering Number of
1 2 3 4 5 6 Classes
@@
Movielens-Text $0.2884_{\pm.02}$ $0.3132_{\pm.09}$ $281_{\pm.07} 0.330$$304_{\pm.05}$ $\mathbf{0.3464_{\pm.08}}$ $0.3217_{\pm.07}$ 5
GoodBook-Text $378_{\pm.06} 0.330$$307_{\pm.03}$ $0.3294_{\pm.03}$ $0.3478_{\pm.05}$ $\mathbf{0.3588_{\pm.04}}$ $0.3227_{\pm.04}$ 4
PPI-Text $0.4919_{\pm.09}$ $0.5371_{\pm.09}$ $0.5581_{\pm.09}$ $\mathbf{0.5613_{\pm.09}}$ $0.5371_{\pm.08}$ $0.5016_{\pm.06}$ 4
Aminer-Text $0.4174_{\pm.02}$ $0.4195_{\pm.02}$ $0.4201_{\pm.02}$ $0.4284_{\pm.02}$ $\mathbf{0.4325_{\pm.02}}$ $0.4278_{\pm.01}$ 3
Cora-CA-Text $0.2535_{\pm.05}$ $0.2645_{\pm.04}$ $0.2637_{\pm.04}$ $0.2716_{\pm.04}$ $0.2633_{\pm.03}$ $\mathbf{0.2685_{\pm.03}}$ 7
Cora-$C$-Text $0.3174_{\pm.03}$ $0.3431_{\pm.04}$ $0.3551_{\pm.04}$ $0.3595_{\pm.05}$ $0.3706_{\pm.04}$ $\mathbf{0.3735_{\pm.04}}$ 7
Pubmed-$C$-Text $0.3531_{\pm.06}$ $0.3577_{\pm.06}$ $0.3856_{\pm.01}$ $0.3918_{\pm.01}$ $\mathbf{0.4009_{\pm.01}}$ $0.3948_{\pm.01}$ 3
Arxiv-Text $0.3345_{\pm.05}$ $0.3315_{\pm.05}$ $0.3198_{\pm.04}$ $0.3329_{\pm.08}$ $\mathbf{0.4189_{\pm.04}}$ $0.3270_{\pm.08}$ 5
IMDB-Text $0.2417_{\pm.01}$ $0.2452_{\pm.01}$ $0.2520_{\pm.04}$ $0.2820_{\pm.05}$ $0.2676_{\pm.04}$ $\mathbf{0.2748_{\pm.04}}$ 5
OGBN-Arxiv $0.2337_{\pm.04}$ $0.2401_{\pm.04}$ $0.2637_{\pm.04}$ $0.2728_{\pm.02}$ $\mathbf{0.2789_{\pm.01}}$ $0.2756_{\pm.04}$ 40
@@ VI Conclusion
In this work, we introduce Hyper-FM, the first Hypergraph Foundation Model for multi-domain knowledge extraction. By implementing Hierarchical High-Order Neighbor Guided Vertex Knowledge Embedding, we integrate domain-specific information into vertex features, addressing varying dimensions and structural complexities. Our Hierarchical Multi-Hypergraph Guided Structural Knowledge Extraction method allows scalable and efficient extraction of structural knowledge from diverse datasets, mitigating negative transfer and domain-specific blending. Additionally, we curate 11 text-attributed hypergraph (TAHG) datasets to bridge hypergraph neural networks and large language models. Empirical evaluations show that Hyper-FM outperforms baseline methods by approximately 10% across these datasets, demonstrating its robustness and effectiveness. We also establish a novel scaling law, revealing that increasing domain diversity significantly enhances model performance, unlike merely expanding vertex and hyperedge counts. Hyper-FM sets a new benchmark for hypergraph foundation models, providing a scalable and effective solution for multi-domain knowledge extraction. Future work will explore expanding TAHG datasets and applying Hyper-FM to a broader range of domains to further enhance its versatility and impact.